OpenStudy (anonymous):

Sixteen more than three-fifths of a number is two less five-ninths of the number. What is the number

OpenStudy (anonymous):

16+3/5=5/9-2 need help?

OpenStudy (anonymous):

is this 9-2 or (5/9)-2?

OpenStudy (anonymous):

(80+3)/5=(5-18)/9 83/5=-13/9

OpenStudy (anonymous):

have u any answer yo this problem?

OpenStudy (anonymous):

Sixteen more than three-fifths of a number is two less five-ninths of the number. What is the number?